# A Generalized Control Revision Method for Autonomous Driving Safety

### v1

@@ III-B Traffic Element Constraint Representation
Following the barrier function (9), based on D-CBF[13] and HOCBF[16] we replace origin CBF constraints with D-CBF constraints and take its derivative. The constraints can be inferred as:
$\begin{split}L_{f}^{2}h+L_{g}L_{f}h\boldsymbol{u}+2\frac{\partial L_{f}h}{\partial\boldsymbol{x}_{ob}}\dot{\boldsymbol{x}}_{ob}+\dot{\boldsymbol{x}}_{ob}^{T}\frac{\partial^{2}h}{\partial\boldsymbol{x}_{ob}^{2}}\dot{\boldsymbol{x}}_{ob}\\ +\frac{\partial\alpha_{1}(h)}{\partial h}\dot{h}+\alpha_{2}(\dot{h}+\alpha_{1}(h))\geqslant 0\end{split}$ (10)
A general selection of extended class $\mathcal{K}_{\infty}$ functions is the linear form, and we will consider $\alpha_{1}$ and $\alpha_{2}$ as linear functions in constraint (10) in this paper later.
Vehicle actuator constrains the applied control input on vehicle system, providing maximum and minimum limits of control input of system (8). Besides, road fiction sets dynamic constraints of vehicle system, and the envelope of the curves of longitudinal force and lateral force is an ellipse, called attachment ellipse. According to these constraints, we calculate control input $\boldsymbol{u^{*}}$ that maximize the barrier function (10), and have the feasibility constraint [17] for integrating CBF:
$\begin{split}h_{F}=L_{f}^{2}h+L_{g}L_{f}h\boldsymbol{u^{*}}+2\frac{\partial L_{f}h}{\partial\boldsymbol{x}_{ob}}\dot{\boldsymbol{x}}_{ob}+\dot{\boldsymbol{x}}_{ob}^{T}\frac{\partial^{2}h}{\partial\boldsymbol{x}_{ob}^{2}}\dot{\boldsymbol{x}}_{ob}\\ +(\alpha_{1}+\alpha_{2})\dot{h}+\alpha_{1}\alpha_{2}h\geqslant 0\end{split}$ (11)
@@
Note that system (8) is an affine system, and it has the same form of control input $\boldsymbol{u}$ (12). This means the constraints of road boundary, described by limiting $d$ in 13, could be combined into the platform discussed above, though two system have different state vectors. Then, the constraint of road boundary can be built following the introduced method.
@@ III-C Integrating CBF
$\underset{\boldsymbol{u}\in\mathbb{U}}{\min}~{}~{}$ $(\boldsymbol{u}-\boldsymbol{u}_{o})^{T}Q(\boldsymbol{u}-\boldsymbol{u}_{o})$ (14a)
s.t. $L_{f}h_{F_{i}}+L_{g}h_{F_{i}}\boldsymbol{u}+\beta_{i}h_{F_{i}}\geqslant 0$ (14b)
$L_{f}h_{r_{j}}+L_{g}h_{r_{j}}\boldsymbol{u}+\gamma_{i}h_{r_{j}}\geqslant 0$ (14c)
where $\boldsymbol{u}_{o}$ is original control input provided by the AD algorithm, and $Q$ is the coefficient matrix to balance the changing of acceleration and steer. Constraint (14b) hopes to avoid potential collision with obstacles, where $i=0,1,..,n-1$ and $n$ is the number of risky obstacles. Constraint (14c) limits vehicle behavior crossing road boundaries, where $j=0,1$.
@@ IV-B4 Ablation study of perception data conversion
To further illustrate the effect of the proposed perception data conversion, we compare the control revision module with and without the supplementary bounding boxes from occupancy grid map on the OnSite simulator platform. As demonstrated in dFig. 9, the black rectangles with red frame lines are the vanilla vectorized perception inputs, and the green rectangle with red frame lines are the supplementary bounding box derived from the process introduced in III-A, which is an object extending out of the truck body. After applying the supplementary perception, the revision module could avoid collision with the irregular obstacle which wasn’t covered by vanilla perception. By adopting both vectorized bounding box and occupancy grid map, our proposed revision method becomes more complete and comprehensive, further guaranteeing driving control safety.

### v2

@@ III-B Traffic Element Constraint Representation
Following the barrier function (9), based on D-CBF[13] and HOCBF[16] we replace origin CBF constraints with D-CBF constraints and take its derivative. Selecting linear functions as extended class $\mathcal{K}_{\infty}$ functions, the constraints can be inferred as:
$\begin{split}&B(\boldsymbol{x},\boldsymbol{u}|h)=L_{f}^{2}h+L_{g}L_{f}h\boldsymbol{u}+2\frac{\partial L_{f}h}{\partial\boldsymbol{x}_{ob}}\dot{\boldsymbol{x}}_{ob}\\ &~{}~{}~{}~{}~{}~{}~{}+\dot{\boldsymbol{x}}_{ob}^{T}\frac{\partial^{2}h}{\partial\boldsymbol{x}_{ob}^{2}}\dot{\boldsymbol{x}}_{ob}+(\alpha_{1}+\alpha_{2})\dot{h}+\alpha_{1}\alpha_{2}h\geqslant 0\end{split}$ (10)
Vehicle actuator constrains the applied control input on vehicle system, providing maximum and minimum limits of control input of system (8). Besides, road fiction sets dynamic constraints of vehicle system, and the envelope of the curves of longitudinal force and lateral force is an ellipse, called attachment ellipse. According to these constraints, we calculate control input $\boldsymbol{u^{*}}$ that maximize the barrier function $B(\boldsymbol{x},\boldsymbol{u})$ (10), and have the feasibility constraint [17] for integrating CBF:
$\begin{split}h_{F}=\underset{\boldsymbol{u^{*}}\in\mathbb{U}}{\max}B(\boldsymbol{x},\boldsymbol{u^{*}}|h)\geqslant 0\end{split}$ (11)
@@
Note that system (12) is an affine system, and it has the same form of control input $\boldsymbol{u}$ in (8). This means constraints of road boundary, described by limiting $d$ in (13), could be combined into the platform discussed above, though two system have different state vectors. Then, the constraint of road boundary can be built following the introduced method.
@@ III-C Integrating CBF
$\underset{\boldsymbol{u}\in\mathbb{U}}{\min}~{}~{}$ $~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}(\boldsymbol{u}-\boldsymbol{u}_{o})^{T}Q(\boldsymbol{u}-\boldsymbol{u}_{o})$ (14a)
s.t. $~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}B(\boldsymbol{x},\boldsymbol{u}|h_{i})\geqslant 0$ (14b)
$~{}~{}L_{f}h_{F_{i}}+L_{g}h_{F_{i}}\boldsymbol{u}+\frac{\partial h_{F_{i}}}{\partial\boldsymbol{x}_{ob}}\dot{\boldsymbol{x}}_{ob}+\beta_{i}h_{F_{i}}\geqslant 0$ (14c)
$L_{f_{r}}^{2}h_{r_{j}}+L_{f_{r}}L_{g_{r}}h_{r_{j}}\boldsymbol{u}+2\gamma_{j}\dot{h}_{r_{j}}+\gamma_{j}^{2}h_{r_{j}}\geqslant 0$ (14d)
where $\boldsymbol{u}_{o}$ is original control input provided by the AD algorithm, and $Q$ is the coefficient matrix to balance the changing of acceleration and steer. Constraint (14b) and (14c) hopes to avoid potential collision with obstacles, where $i=0,1,..,n-1$ and $n$ is the number of risky obstacles. Constraint (14d) limits vehicle behavior crossing road boundaries, where $j=0,1$ and $f_{r},g_{r}$ are from system (12).
@@ IV-B4 Ablation study of perception data conversion
To further illustrate the effect of the proposed perception data conversion, we compare the control revision module with and without the supplementary bounding boxes from occupancy grid map on the OnSite simulator platform. As demonstrated in Fig. 9, the black rectangles with red frame lines are the vanilla vectorized perception inputs, and the green rectangle with red frame lines are the supplementary bounding box derived from the process introduced in III-A, which is an object extending out of the truck body. After applying the supplementary perception, the revision module could avoid collision with the irregular obstacle which wasn’t covered by vanilla perception. By adopting both vectorized bounding box and occupancy grid map, our proposed revision method becomes more complete and comprehensive, further guaranteeing driving control safety.